# MR. Guard: Multilingual Reasoning Guardrail using Curriculum Learning

### v1

@@ Abstract
Large Language Models (LLMs) are susceptible to adversarial attacks such as jailbreaking, which can elicit harmful or unsafe behaviors. This vulnerability is exacerbated in multilingual setting, where multilingual safety-aligned data are often limited. Thus, developing a guardrail capable of detecting and filtering unsafe content across diverse languages is critical for deploying LLMs in real-world applications. In this work, we propose an approach to build a multilingual guardrail with reasoning. Our method consists of: (1) synthetic multilingual data generation incorporating culturally and linguistically nuanced variants, (2) supervised fine-tuning, and (3) a curriculum-guided Group Relative Policy Optimization (GRPO) framework that further improves performance. Experimental results demonstrate that our multilingual guardrail consistently outperforms recent baselines across both in-domain and out-of-domain languages. The multilingual reasoning capability of our guardrail enables it to generate multilingual explanations, which are particularly useful for understanding language-specific risks and ambiguities in multilingual content moderation.
@@
MR. Guard: Multilingual Reasoning Guardrail using Curriculum Learning
@@ 1 Introduction
Figure 1: The guardrails are providing different predictions for English and Chinese inputs with the same semantic meaning. In this case, we are using LLaMa-Guard-3 as the guardrail model.
To address these issues, safety alignment techniques—such as Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO)—have been proposed to align LLM behavior with human values and reduce unsafe generations (Ouyang et al., 2022; Rafailov et al., 2023). Another line of work focuses on building standalone safety classifiers or guardrails, which act as filters to detect and block unsafe user prompts or model generations without modifying the LLM itself (Inan et al., 2023; Ghosh et al., 2024). These lightweight safety modules are advantageous in being more efficient and easier to deploy or update (Table 1). Most existing methods are primarily English-centric (Liu et al., 2025; Kang and Li, 2024b; Yuan et al., 2024) which cannot handle multilingual content moderation (Yang et al., 2024). As shown in Figure 1, the guardrail model successfully identifies the unsafe user prompt in English but fails to detect its semantically equivalent counterpart in Chinese. Moreover, without explanations, it becomes difficult to understand the rationale behind the guardrail’s decisions [^1] We interchangeably use guard and guardrail through the paper.
@@
  • We propose an approach which leverages curriculum learning (Bengio et al., 2009a) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024b) to train a multilingual guardrail model with reasoning ability and show that it can distinguish safe/unsafe user prompt across various languages.
  • We show state-of-the-art performance of our proposed multilingual guardrail on various multilingual safety benchmark over multiple baselines in performance.
  • We evaluate the robustness of multilingual guardrails on multilingual variations such as code-switching and sandwich attack.
@@ 2.1 Multilingual LLM Safety
While LLMs demonstrate strong cross-lingual capabilities on multilingual downstream tasks, their ability to handle unsafe content in multilingual settings remains largely unknown, and there is still significant room for improving their robustness to multilingual inputs. Prior studies (Wang et al., 2023; Deng et al., 2023) have shown that LLMs are vulnerable to non-English jailbreaking prompts, especially in low-resource languages. Follow-up work (Yoo et al., 2024) uses GPT-4 to combine parallel jailbreaking queries in Deng et al. (2023) from different languages into a single code-switching prompt, demonstrating that such prompts further increase the attack success rate compared to monolingual attacks. Recent work (de Wynter et al., 2024; Jain et al., 2024; Ye et al., 2023) collects multilingual moderation datasets to investigate the ability of LLM to respond to multilingual harmful prompts/responses and assess whether guardrails effectively filter out them. They all show that existing guard or encoder-only classifiers cannot adequately handle multilingual content moderation. Upadhayay and Behzadan (2024) has introduced an attack against LLMs by embedding jailbreaking prompts within unrelated, low-resource, but safe inputs. Their results show that both proprietary and open-source models are vulnerable to these low-resource distractors, often following the embedded unsafe instructions and generating harmful content. These findings all underscore the urgent need for robust guardrails capable of detecting and mitigating unsafe behavior in multilingual settings.
@@ 2.2 Guardrails for safeguarding LLMs
Recent guardrail research (Li et al., 2024; Inan et al., 2023; Rebedea et al., 2023; Ghosh et al., 2025; Kang and Li, 2024a) has leveraged pre-trained small language models (SLMs), such as LlaMa-2/3.1-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023), to distinguish between safe and unsafe content. These methods have demonstrated promising results in detecting harmful inputs in English compared to encoder-only models. Yuan et al. (2024) has enhanced base guardrail models by incorporating energy-based data generation and combining guardrail predictions with k-nearest neighbors (kNN) predictions. Additionally, Kang and Li (2024b) has introduced a knowledge-based logical reasoning framework, which first ask the model to determine whether an input belongs to a predefined risk category and then uses a probabilistic graphical model to estimate the likelihood of unsafety. GuardReasoner (Liu et al., 2025) has improved interpretability and performance by training the base model on reasoning-augmented data and applying reinforcement learning (RL) algorithm DPO (Rafailov et al., 2023) to select difficult examples. However, these efforts largely focus on English. To address multilingual safety, Deng et al. (2025) has introduced an RL-based method for generating synthetic multilingual data by iteratively and jointly updating a synthetic data generator and a guardrail model. Yet, their work targets only high-resource languages close to English. In contrast, we focus on building a multilingual guardrail capable of handling cultural and language-specific challenges, including languages from different language families.
To enhance the reasoning capabilities of language models, we integrate curriculum learning with a reinforcement learning (RL) algorithm known as Group Relative Policy Optimization (GRPO). GRPO demonstrates superior performance compared to offline methods such as Direct Preference Optimization (DPO)(Rafailov et al., 2023), while offering improved computational efficiency over on-policy algorithms like Proximal Policy Optimization (PPO). A more comprehensive discussion of these RL approaches and their relationship to curriculum learning is provided in Appendix A.
@@ 3 Multilingual Guard with reasoning
We detail our algorithm for building a guardrail with reasoning capabilities for multilingual content moderation in this section. The approach consists of three key components as shown in Figure 2:
1. Synthetic Data Generation – Inspired by Liu et al. (2025), we collect the analysis of safety for our seed data from a more powerful proprietary model, GPT-4o-mini [^2] We utilize GPT-4o-mini (gpt-4o-mini-2024-07-18) through all data generation stage. and use the collected data to train our model. We additionally generate multilingual data and corresponding multilingual analysis using GPT-4o-mini.
@@
3. Curriculum-Based Optimization – We combine a three-stage curriculum learning framework and Group Relative Policy Optimization (GRPO) to align the model with desired multilingual moderation behavior.
@@ 3.1 Synthetic Data Generation
Next, we subsample a smaller set from the original dataset $D$, forming a subset $D_{\text{sub}}$. For each target language $l_{k}$, we prompt GPT to translate each English prompt $p_{i}^{l_{0}}\in D_{\text{sub}}$ into the target language, resulting in $p_{i}^{l_{k}}$. We assume that the safety label $y_{i}$ is preserved across translations. To further ensure label consistency, we prompt GPT to reassess the safety of each translated prompt $p_{i}^{l_{k}}$, If the reassessed label conflicts with the original $y_{i}$, the corresponding example is discarded from the training set. To encourage the diversity of reasoning across languages, we generated both English reasoning $e^{l_{0}}_{i}$ and reasoning in the original language $e^{l_{k}}_{i}$. We then prompt GPT with $(p_{i}^{l_{k}},y_{i})$ to generate the corresponding reasoning $e_{i}^{l_{k}}$ in language $l_{k}$ and English $e_{i}^{l_{0}}$, yielding the dataset $D^{l_{k}}=\{(p_{i}^{l_{k}},e_{i}^{l_{0}},e_{i}^{l_{k}},y_{i})\}_{i=1}^{n}.$. Note that $D^{l_{k}}$ contains significantly fewer examples than $D^{l_{0}}$. Given $K$ target languages $\{l_{1},\ldots,l_{K}\}$, we obtain the multilingual dataset:
@@ 3.2 Supervised Fine-tuning
We perform supervised fine-tuning of the base model, denoted by $\pi$, using the multilingual dataset $D^{\text{multi}}$, to enable the model to identify safe and unsafe prompts along with their reasoning. Given a data point $(p_{i}^{l_{k}},e_{i}^{l_{0}},e_{i}^{l_{k}},y_{i})$, we fine-tune the base model by applying gradient descent on the tokens corresponding to both reasoning trajectories and the safety label. This enables the model to leverage the strong generalization capabilities of English while simultaneously developing multilingual reasoning skills, thereby preparing it for the subsequent reinforcement learning stage. We denote the resulting fine-tuned model as $\pi_{\text{ref}}$.
3.3 Curriculum-guided GRPO
In this stage, we employ reinforcement learning to further enhance detection performance by eliciting stronger reasoning capabilities. We begin by re-sampling a subset $D^{l_{0}^{\prime}}$ from the original English safety training dataset $D^{l_{0}}$. Each prompt in $D^{l_{0}^{\prime}}$ is then translated into the target languages $l_{k}$, for $k\in\{1,\ldots,K\}$. [^3] The resulting number of valid multilingual data points may be smaller than $n^{\prime}$ due to potential GPT translation refusals or changes in safety labels. We then introduce a curriculum-based training schedule. The intuition is that, since the base model is initially fine-tuned on an English-dominant corpus, it is more familiar with English-specific nuances, such as slang and native expressions. To guide the model in progressively learning to handle other languages as second languages, we propose a curriculum that gradually introduces more challenging native multilingual variants. These variants are derived from English sentences and are incorporated stage by stage to support step-wise multilingual adaptation. To construct the curriculum, we introduce a novel difficulty function $\operatorname{Diff}$ that quantifies the difficulty of prompts in the target languages. Specifically, all English prompts are assigned a baseline difficulty level of 0. For a prompt $p^{l_{k}}$ in language $l_{k}$ and its corresponding English prompt $p^{l_{0}}\in D^{l_{0}^{\prime}}$, we use the prompt template shown in Figure 6 to instruct GPT to generate two challenging variants, $p^{l_{k}\prime}$ and $p^{l_{k}\prime\prime}$, enriched with slang, references to local places, institutions, foods, and other culturally or linguistically specific elements. A translation model $\pi_{\text{bt}}$ is then used to back-translate $p^{l_{k}\prime}$ and $p^{l_{k}\prime\prime}$ into English. The semantic similarity between the back-translated prompt and the original English prompt $p^{l_{0}}$ is computed using cosine distance function $\operatorname{cos}$. The difficulty level of a back-translated prompt $p\in\{p^{l_{k}},p^{l_{k}\prime},p^{l_{k}\prime\prime}\}$is defined as:
@@ 3.3 Curriculum-guided GRPO
After developing the curriculum, we apply GRPO to optimize the reference model $\pi_{\text{ref}}$ (Shao et al., 2024b). We utilize rule-based reward functions, with the following components:
Format reward ($\mathcal{R}_{f}$)
: This reward penalizes formatting errors. If the output does not contain a properly formatted safety prediction (i.e., "Safety: safe" or "Safety: unsafe") which is often happened in multilingual generation, the reward is -1. Otherwise, the reward is 1.
Correctness reward ($\mathcal{R}_{c}$)
: If the safety prediction is correct, the reward is 1, otherwise, the reward is -1.
Uncertainty reward ($\mathcal{R}_{u}$)
: We train an auxillary encoder-only model $\pi_{u}$ to use the reasoning to decide whether the input is safe or not (binary classification) and take the softmax score as reward.
@@
Language reward ($\mathcal{R}_{lang}$)
: For the second and third stages, the input sentences are more native to the target language. We hypothesize that language-specific reasoning enhances the model’s understanding in these settings. To encourage the model to generate reasoning in the target language, we add this language reward,
@@
With the reward signals defined, we apply the original GRPO algorithm, whose objective function is described as follows. Let $\pi_{\theta}$ denote the language model parameterized by $\theta$, and $\pi_{\text{old}}$ be the model from a previous iteration. Given a prompt $p$, we sample a group of generations $\{o_{1},\ldots,o_{G}\}\sim\pi_{\text{old}}(\cdot\mid p)$, each associated with a reward $\{r_{1},\ldots,r_{G}\}$. Let $\mathbb{D}_{\text{KL}}$ denote the KL-divergence between two distributions. GRPO estimates the advantage of a generation $o_{i}$ using:
The GRPO objective is then defined as:
$-\beta\mathbb{D}_{\rm KL}[\pi_{\theta}||\pi_{\rm ref}].$
For a detailed explanation of GRPO, please refer to Shao et al. (2024a).
@@ 4.1 Experimental Setup
In our experiments, we use the training set from Aegis-2.0-Safety (Ghosh et al., 2025) as the English seed data. Our base model is LLaMA-3.1-8B-Instruct, and we apply QLoRA (Dettmers et al., 2023) for parameter-efficient fine-tuning during both the SFT and GRPO stage [^4] More details on datasets and training are in B. To construct the curriculum using back-translation, we employ the facebook/nllb-200-3.3B model for translation and use all-MiniLM-L6-v2 to compute sentence embeddings of both the original and back-translated sentences. For the difficulty threshold, we set $t_{1}=0.85$ and $t_{2}=0.7$. To determine the language of the sampled output, we utilize papluca/xlm-roberta-base-language-detection. Our experiments divide the test sets into two categories: in-domain languages, which are covered during training, and out-of-domain languages, which are not seen during training.
@@
SFT 85.78 81.15 95.74 94.14 90.44 86.02 86.12 86.88 94.97 92.09
SFT+GRPO 90.35 83.82 97.33 97.02 92.69 89.66 89.64 90.21 96.76 93.22
SFT+Curr+GRPO 91.08 85.87 98.16 98.21 94.33 92.06 91.17 92.15 97.26 95.74
Table 2: Performance of different guardrails to identify multilingual safety across five benchmark datasets. We report F1 scores as the evaluation metric and bold the best-performing results for each dataset where ID refers to in-domain languages and OOD refers to out-of-domain languages. Top are baselines and the bottom part are our model.

### v2

# MrGuard: A Multilingual Reasoning Guardrail for Universal LLM Safety

@@ Abstract
Large Language Models (LLMs) are susceptible to adversarial attacks such as jailbreaking, which can elicit harmful or unsafe behaviors. This vulnerability is exacerbated in multilingual settings, where multilingual safety-aligned data is often limited. Thus, developing a guardrail capable of detecting and filtering unsafe content across diverse languages is critical for deploying LLMs in real-world applications. In this work, we introduce a multilingual guardrail with reasoning for prompt classification. Our method consists of: (1) synthetic multilingual data generation incorporating culturally and linguistically nuanced variants, (2) supervised fine-tuning, and (3) a curriculum-based Group Relative Policy Optimization (GRPO) framework that further improves performance. Experimental results demonstrate that our multilingual guardrail, MrGuard, consistently outperforms recent baselines across both in-domain and out-of-domain languages by more than 15%. We also evaluate MrGuard’s robustness to multilingual variations, such as code-switching and low-resource language distractors in the prompt, and demonstrate that it preserves safety judgments under these challenging conditions. The multilingual reasoning capability of our guardrail enables it to generate explanations, which are particularly useful for understanding language-specific risks and ambiguities in multilingual content moderation.
@@
MrGuard: A Multilingual Reasoning Guardrail for Universal LLM Safety
@@ 1 Introduction
Figure 1: The guardrails specialized in English (GuardR, (Liu et al., 2025)) are providing different predictions for English and Chinese inputs with the same semantic meaning. Our MrGuard can analyze the Chinese prompts with explanation and provide correct safety prediction.
@@
To address these issues, safety alignment techniques—such as Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO)—have been proposed to align LLM behavior with human values and reduce unsafe generations (Ouyang et al., 2022; Rafailov et al., 2023). Another line of work focuses on building standalone safety classifiers or guardrails, which act as filters to detect and block unsafe user prompts or model generations without modifying the LLM itself (Inan et al., 2023; Ghosh et al., 2024). These lightweight safety modules are advantageous in being more efficient and easier to deploy or update (Table 1). Most existing methods are primarily English-centric (Liu et al., 2025; Kang and Li, 2024b; Yuan et al., 2024) which cannot handle multilingual content moderation (Yang et al., 2024). As shown in Figure 1, the guardrail model successfully identifies the unsafe user prompt in English but fails to detect its semantically equivalent counterpart in Chinese. Moreover, without explanations, it becomes difficult to understand the rationale behind the guardrail’s decisions [^1] We interchangeably use guard and guardrail through the paper..
@@
  • We introduce MrGuard, a multilingual reasoning-enhanced guardrail for prompt moderation that improves performance and robustness across languages. Our approach combines curriculum learning (Bengio et al., 2009a) with Group Relative Policy Optimization (GRPO) (Shao et al., 2024b) to gradually introduce more culturally diverse variants at the post-training stage.
  • We achieve state-of-the-art results on several multilingual safety benchmarks, outperforming all baselines in prompt classification accuracy. We further demonstrate MR.Guard’s robustness on multilingual variations such as code-switching and sandwich Upadhayay and Behzadan (2024) attacks. Our results show that post-training with reasoning abilities significantly improves the robustness and performance of guardrails on multilingual prompt classification.
  • We present a comprehensive evaluation of our reasoning-enhanced guardrail with key metrics such as cross-lingual consistency and reasoning fidelity which establishes a strong baseline for future assessments of guardrail reasoning capabilities.
@@ 2.1 Multilingual LLM Safety
While LLMs demonstrate strong cross-lingual capabilities on multilingual downstream tasks, their ability to handle unsafe content in multilingual settings remains largely unknown, and there is still significant room for improving their robustness to multilingual inputs. Prior studies (Wang et al., 2023; Deng et al., 2023) have shown that LLMs are vulnerable to non-English jailbreaking prompts, especially in low-resource languages. Follow-up work (Yoo et al., 2024) uses GPT-4 to combine parallel jailbreaking queries in Deng et al. (2023) from different languages into a single code-switching prompt, demonstrating that such prompts further increase the attack success rate compared to monolingual attacks. Recent work (de Wynter et al., 2024; Jain et al., 2024; Ye et al., 2023) collects multilingual moderation datasets to investigate the ability of LLMs to respond to multilingual harmful prompts/responses and assess whether guardrails effectively filter them out. They all show that existing guard or encoder-only classifiers cannot adequately handle multilingual content moderation. Upadhayay and Behzadan (2024) has introduced an attack against LLMs by embedding jailbreaking prompts within unrelated, low-resource, but safe inputs. Their results show that both proprietary and open-source models are vulnerable to these low-resource distractors, often following the embedded unsafe instructions and generating harmful content. These findings all underscore the urgent need for robust guardrails capable of detecting and mitigating unsafe behavior in multilingual settings.
@@ 2.2 Guardrails for safeguarding LLMs
Recent guardrail research (Li et al., 2024; Inan et al., 2023; Rebedea et al., 2023; Ghosh et al., 2025; Kang and Li, 2024a) have leveraged pre-trained small language models (SLMs), such as LlaMa-2/3.1-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023), to distinguish between safe and unsafe content. These methods have demonstrated promising results in detecting harmful inputs in English compared to encoder-only models. Yuan et al. (2024) has enhanced base guardrail models by incorporating energy-based data generation and combining guardrail predictions with k-nearest neighbors (kNN) predictions. Additionally, Kang and Li (2024b) has introduced a knowledge-based logical reasoning framework, which first asks the model to determine whether an input belongs to a predefined risk category and then uses a probabilistic graphical model to estimate the likelihood of unsafety. GuardReasoner (Liu et al., 2025) has improved interpretability and performance by training the base model on reasoning-augmented data and applying reinforcement learning (RL) algorithm DPO (Rafailov et al., 2023) to select difficult examples. However, these efforts largely focus on English. To address multilingual safety, Deng et al. (2025) has introduced an RL-based method for generating synthetic multilingual data by iteratively and jointly updating a synthetic data generator and a guardrail model. Yet, their work targets only high-resource languages close to English. In contrast, we focus on building MrGuard: a multilingual guardrail capable of handling cultural nuances and language-specific challenges spanning languages from several different families.
To enhance the reasoning capabilities of language models, we integrate curriculum learning with a reinforcement learning (RL) algorithm known as Group Relative Policy Optimization (GRPO) Shao et al. (2024b). GRPO demonstrates superior performance compared to offline methods such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), while offering improved computational efficiency over on-policy algorithms like Proximal Policy Optimization (PPO). A more comprehensive discussion of these RL approaches and their relationship to curriculum learning is provided in Appendix A.
@@ 3 Multilingual Guard with Reasoning
We detail our algorithm for building MrGuard: a guardrail with reasoning capabilities for multilingual content moderation in this section. The approach consists of three key components as shown in Figure 2:
1. Synthetic Data Generation – Inspired by Liu et al. (2025), we collect the analysis of safety for our seed data from a more powerful proprietary model, GPT-4o-mini [^2] We utilize GPT-4o-mini (gpt-4o-mini-2024-07-18) through all data generation stage, referred to as GPT for simplcity, in the next subsection. and use the collected data to train our model. We additionally generate multilingual data and corresponding multilingual analysis using GPT-4o-mini.
@@
3. Curriculum-Based Optimization – We combine a three-stage curriculum learning framework and GRPO Shao et al. (2024b) to align the model with desired multilingual moderation behavior.
@@ 3.1 Synthetic Data Generation
Next, we subsample a smaller set from the original dataset $D$, forming a subset $D_{\text{sub}}$. For each target language $l_{k}$, we prompt GPT to translate each English prompt $p_{i}^{l_{0}}\in D_{\text{sub}}$ into the target language, resulting in $p_{i}^{l_{k}}$. We assume that the safety label $y_{i}$ is preserved across translations. To further ensure label consistency, we prompt GPT to reassess the safety of each translated prompt $p_{i}^{l_{k}}$, If the reassessed label conflicts with the original $y_{i}$, the corresponding example is discarded from the training set. We then prompt GPT with $(p_{i}^{l_{k}},y_{i})$ to generate the corresponding reasoning $e_{i}^{l_{k}}$ in language $l_{k}$ and English $e_{i}^{l_{0}}$, yielding the dataset $D^{l_{k}}=\{(p_{i}^{l_{k}},e_{i}^{l_{0}},e_{i}^{l_{k}},y_{i})\}_{i=1}^{n}.$. Note that $D^{l_{k}}$ contains significantly fewer examples than $D^{l_{0}}$. Given $K$ target languages $\{l_{1},\ldots,l_{K}\}$, we obtain the multilingual dataset:
@@ 3.2 Supervised Fine-tuning
We perform supervised fine-tuning of the base model, denoted by $\pi$, using the multilingual dataset $D^{\text{multi}}$, to enable the model to identify safe and unsafe prompts along with their reasoning. Given a data point $(p_{i}^{l_{k}},e_{i}^{l_{0}},e_{i}^{l_{k}},y_{i})$, we fine-tune the base model by applying cross-entropy loss on the tokens corresponding to both reasoning trajectories and the safety label. This enables the model to leverage the strong generalization capabilities of English while simultaneously developing multilingual reasoning skills, thereby preparing it for the subsequent reinforcement learning stage. We denote the resulting fine-tuned model as $\pi_{\text{sft}}$.
3.3 Curriculum-based GRPO
In this stage, we employ reinforcement learning to further enhance detection performance by eliciting stronger reasoning capabilities. We begin by re-sampling a subset $D^{l_{0}^{\prime}}$ from the original English safety training dataset $D^{l_{0}}$. Each prompt in $D^{l_{0}^{\prime}}$ is then translated into the target languages $l_{k}$, for $k\in\{1,\ldots,K\}$.We then introduce a curriculum-based training schedule. The intuition is that, since the base model is initially fine-tuned on an English-dominant corpus, it is more familiar with English-specific nuances, such as slang and native expressions. To guide the model in progressively learning to handle other languages as second languages, we propose a curriculum that gradually introduces more challenging native multilingual variants. These variants are derived from English sentences and are incorporated stage by stage to support step-wise multilingual adaptation. To construct the curriculum, we introduce a novel difficulty function $\operatorname{Diff}$ that quantifies the difficulty of prompts in various target languages. Specifically, all the English prompts are assigned a baseline difficulty level of 0. For a prompt $p^{l_{k}}$ in language $l_{k}$ and its corresponding English prompt $p^{l_{0}}\in D^{l_{0}^{\prime}}$, we use the prompt template shown in Figure 6 to instruct GPT to generate two challenging variants, $p^{l_{k}\prime}$ and $p^{l_{k}\prime\prime}$, enriched with slang, references to local places, institutions, foods, and other culturally or linguistically specific elements. A translation model $\pi_{\text{bt}}$ is then used to back-translate $p^{l_{k}\prime}$ and $p^{l_{k}\prime\prime}$ into English. The semantic similarity between the back-translated prompt and the original English prompt $p^{l_{0}}$ is computed using the cosine similarity function $\operatorname{cos}$. The difficulty of a back-translated prompt $p\in\{p^{l_{k}},p^{l_{k}\prime},p^{l_{k}\prime\prime}\}$ is defined as:
@@ 3.3 Curriculum-based GRPO
After developing the curriculum, we apply GRPO to optimize the reference model $\pi_{\text{sft}}$ (Shao et al., 2024b). We utilize rule-based reward functions, with the following components:
Format reward ($\mathcal{R}_{f}$):
This reward penalizes formatting errors. If the output does not contain a properly formatted safety prediction (i.e., "Safety: safe" or "Safety: unsafe") which often happens in multilingual generation, the reward is $-1$. Otherwise, the reward is $1$.
Correctness reward ($\mathcal{R}_{c}$):
If the safety prediction is correct, the reward is $1$, otherwise, the reward is $-1$.
Uncertainty reward ($\mathcal{R}_{u}$):
We train an auxiliary encoder-only model $\pi_{u}$ to use the reasoning to decide whether the input is safe or not (binary classification) and take the softmax score as the reward.
@@
Language reward ($\mathcal{R}_{lang}$):
For the second and third stages, the input sentences are more native to the target language. We hypothesize that language-specific reasoning enhances the model’s understanding in this setting. To encourage the model to generate reasoning in the target language, we add this language reward,
@@
With the reward signals defined, we apply the original GRPO algorithm to optimize the reference model $\pi_{\text{sft}}$. For a detailed explanation of GRPO, please refer to Appendix E and Shao et al. (2024a).
@@ 4.1 Experimental Setup
In our experiments, we use the training set from Aegis-2.0-Safety (Ghosh et al., 2025) as the English seed data. Our base model is LLaMA-3.1-8B-Instruct (Aaron Grattafiori, 2024), and we apply QLoRA (Dettmers et al., 2023) for parameter-efficient fine-tuning during both the SFT and GRPO stage [^3] More details on the training hyperparameters and configurations are in Appendix B.. To construct the curriculum using back-translation, we employ the facebook/nllb-200-3.3B model for translation and use all-MiniLM-L6-v2 to compute the sentence embeddings of both the original and the back-translated sentences. For the difficulty threshold, we set $t_{1}=0.85$ and $t_{2}=0.7$. To determine the language of the sampled output, we utilize an xlm-based language detector  [^4] The full name of the language detector: papluca/xlm-roberta-base-language-detection. Our experiments divide the test sets into two categories: in-domain languages, which are covered during training, and out-of-domain languages, which are not seen during training.
@@
MrGuard (Ours) 90.95 86.90 98.18 98.11 94.56 92.06 91.45 92.68 97.41 96.15
Table 2: Performance of different guardrails to identify multilingual safety across five benchmark datasets. We report F1 scores as the evaluation metric and bold the best-performing results for each dataset where ID refers to in-domain languages and OOD refers to out-of-domain languages. Top are baselines and the bottom part is MrGuard (ours). The model size and training dataset size are listed in Table 1.